OpenStudy (anonymous):

Can someone help me with this one? Find the indicated outputs for f(x)=3X2-2X F(0)= F(-1)= F(2)=

OpenStudy (saifoo.khan):

insert 0, -1 & 2 one by one instead of x.

OpenStudy (anonymous):

Like Saifoo said. Just substitute the value for x and solve the equation. Here's an example. Suppose G(x) = 2x Well, to find G(1), I just substitute 1 in for x and solve, so G(1) = 2(1) = 2. What's G(5)? G(5) = 2(5) = 10 G(8) = 2(8) = 16 And so on.

OpenStudy (saifoo.khan):

hmm.

OpenStudy (anonymous):

Whatcha thinkin'?

OpenStudy (anonymous):

Im really thinking of blowing my goddamn head off Im so confused

OpenStudy (saifoo.khan):

lol, there is nothing to get confused.

OpenStudy (anonymous):

Haha okay, well let's work through this so you aren't confused anymore. What's tricky about it to you?

OpenStudy (anonymous):

This can be real hard to understand at first, but I'm sure that you can get it.

OpenStudy (anonymous):

I dont even know what an indicated output is

OpenStudy (anonymous):

When they say indicated output, they just mean the result. Basically a function takes in an input, or an x, and gives out an output, often called a y or F(x). So F(x) means that F is the function, and x is the input. F(x) is the output for any given x. So F(2) means, "the output of the function when the input is 2."

OpenStudy (anonymous):

can you explain a little better?

OpenStudy (anonymous):

Look a juicer is a device to prepare juice if u put apple u get apple juice, if mango than mango and so forth . Similarly consider a function to b a juicer if u put one number it results one output if another then another. thats it

OpenStudy (anonymous):

If u put 0 u get f(0)=3(0)^2-2(0)=0-0=0 If u put -1 u get f(-1)=3(-1)^2-2(-1)=3+2=6 If u put 2u get f(2)=3(2)^2-2(2)=12-4=8 that what it says

OpenStudy (anonymous):

???

OpenStudy (anonymous):

u r really a kid. do some practice

OpenStudy (anonymous):

Learn some people skills instead of being rude

OpenStudy (anonymous):

m not rude. I m just giving u the direction